OpenStudy (anonymous):

find the inverse LaPlace transform of the following function: F(s) = (e^(-s) - e^(-4s))/ s^2

OpenStudy (anonymous):

\[F(s) = (e ^{-s} - e ^{-4s})/ s ^{2}\]

OpenStudy (turingtest):

F(s) = (e^(-s) - e^(-4s))/ s^2\[F(s)=e^{-s}/s^2-e^{-4s}/s^2\] we're going to do this as a heaviside function right? just thinking how

OpenStudy (jamesj):

The Heaviside function/unit step function u(x) has Laplace transform 1/s. The function u(x-a) has Laplace transform e^(-as)/s So now you should have it.

OpenStudy (jamesj):

Just be sure to simplify your answer. I.e., you can simplify and get rid of all the u(.) in the final answer; it is instructive if you do.

OpenStudy (anonymous):

How do you get rid of the u in the final answer?

OpenStudy (jamesj):

Oh sorry. You have 1/s^2. The Laplace transform of t is 1/s^2

OpenStudy (jamesj):

Now the e^(-as) are shift functions

OpenStudy (turingtest):

\[F(s)=e^{-s}(1/s^2)-e^{-4s}(1/s^2)\]yeah so we have do do it like e^(-s)G(s)

OpenStudy (jamesj):

That is, if a function f(t) has Laplace transform F(s), then the function f(t-a) has Laplace transform e^(-as)F(s)

OpenStudy (jamesj):

Hence, for example, the inverse transform of e^(-1s)/s^2 is (t-1)

OpenStudy (anonymous):

so I would re write my equation as a step function and then take the LaPlace transform?

OpenStudy (jamesj):

No -- now that realize we have 1/s^2 terms, we have no step functions involved.

OpenStudy (jamesj):

Look at what I just did with e^(-1s)/s^2

OpenStudy (anonymous):

sorry I meant to say shift function

OpenStudy (anonymous):

the answer is just (t-1) - (t-4)

OpenStudy (jamesj):

Now be careful about domains.

OpenStudy (jamesj):

The inverse Laplace transform of e^(-s)/s^2 is t - 1, for t >= 1 and 0 elsewhere

OpenStudy (anonymous):

therefore I need to include the Uc(t) function

OpenStudy (jamesj):

Likewise the inverse Laplace transform of e^(-4s)/s^2 is t - 4, t >=4 0, t < 4 So when you add these two functions you get what?

OpenStudy (jamesj):

or subtract them actually.

OpenStudy (anonymous):

Ok thank you, I do want to use the Uc(t) expression then.

OpenStudy (jamesj):

The final answer is f(t) = 0 , t < 1 = t - 1 , t in [1,4] = (t-1) - (t-4), t > 4 Now simplify this.

OpenStudy (anonymous):

I understand my instructor wants us to use Uc(t) expression he taught us, which gives the domain. Thank you so much

OpenStudy (jamesj):

ok

OpenStudy (anonymous):

I have another question. Find the LaPlace transform of \[f(t) = 1/\sqrt{t}\]

OpenStudy (turingtest):

I can't do that one, but it's number 4 on this chart: http://tutorial.math.lamar.edu/pdf/Laplace_Table.pdf I'm not sure how to use that capital Gamma thing though. I'm sure JamesJ does. You may need to repost the question.

OpenStudy (anonymous):

thank you

OpenStudy (jamesj):

Just saw this. Look up the gamma function; you'll find that the Laplace (not LaPlace ;-) ) transform is just applying the definition of that function to 1/sqrt(t) and the fact of the matter is except for integral value of gamma function, it's not meaningfully reducible to elementary functions.

OpenStudy (turingtest):

blech, sounds ugly. I've heard of the Gamma function. Better study up on it.

OpenStudy (jamesj):

I'm sure we all thought sin and cos sounded a bit complicated at first, but we later figured out they're quite elegant. Something is true with the Gamme Function too. It's quite nice.